Malma is a HK manufacturer that has recently changed suppliers and has started importing one of its key components from Germany. The Finance Director has suggested that the company hedge all future Euro foreign currency transactions however the Managing Director thinks that the Euro is very stable and that this is a cost that could be avoided.

The first shipment has arrived and Malma are due to pay 2,700,000 Euros in 3 months’ time. The following information has been provided in order to assess the use of a forward contract and a money market hedge:

                                           

Current spot exchange rate       0.1285 – 0.1290 Euro / HK$

3 month future exchange rate    0.1288 – 0.1294 Euro / HK$

Annual Euro borrowing rate       6% per annum

Annual Euro deposit rate            3% per annum

Annual HK deposit rate               4% per annum

Annual HK borrowing rate          7% per annum

Malma currently has a cash surplus.

Discuss how exchange rates move if interest rates in the two countries are different and conclude whether this would support the Managing Director’s view of not hedging the transaction. (You should use the average figures above to illustrate your answer)
Answer #1

As per Interest Rate Parity (IRP), the size of forward premium (or discount) should be equal to the interest rate differential between the two countries of concern, i.e., differential interest rate between countries effect forward exchange rate to premium (or discount).

The Formula for Covered Interest Rate Parity Is

(1+i_{d})=\frac{F}{S}*(1+i_{f})

where:

id​=The interest rate in the domestic currency or the base currency

if= The interest rate in the foreign currency or the quoted currency

S=The current spot exchange rate

F=The forward foreign exchange rate​

Using the above formula calculating the 3-month forward exchange rate

{1+ .04*(3/12)} = (F/.1285)*(1+.03*(3/12))

1.01 = (F/.1285)*1.0075

F = 0.1288 Euro/HK$

Since, as per Interest rate parity, with interest rate in Hong Kong and Euro the exchange rate in 3-months will be 0.1288 Euro/HK$ which is same as of 3-month Forward Exchange Rate.

Hence, 3-month Forward rate and expected exchange rate in 3 months is same Managing Director's view of Euro being stable is right and Malma need not to Hedge its position in Forward Contract.

Now, Analysing Managing Director's point regarding Money Market Hedge

So, in Money market hedge, we will first Borrow HK$ ( being opportunity cost lost), then convert it into Euro, & then invest Euro and then repay Euro in 3 months.

Amount in Euro to be invested to repay EURO in 3 months= 2700,000/(1+(.03*3/12))

= Euro 2679900

Amount of HK$ needed to convert to Euro = Euro 2679900/0.1285

= Hk$ 20,855,253

So, Malma will needed HK$ 20,855,253 now to convert it into Euro and then to invest it.

Interest cost of borrowing HK$ (or Opportunity Cost) = 20855253*(0.07*3/12)

= Hk$ 364,967

So, total amount to be paid after 3-months as per Money Market Hedge= HK$ 21,220,220

If, EURO 2700,000 is paid as per expected 3-month future rate of 0.1288 Euro/HK$

then total amount to be paid after 3- months = 2700000/0.1288

= HK$ 20,962,733

​Money Market Hedge option is also not viable as less HK$ is to be paid without Hedging.

So, managing Director's view is right and Malma should not hedge its foreign Currancy.